# Resource-Constrained Federated Continual Learning: What Does Matter?

### v1

@@ Abstract
Federated Continual Learning (FCL) aims to enable sequentially privacy-preserving model training on streams of incoming data that vary in edge devices by preserving previous knowledge while adapting to new data. Current FCL literature focuses on restricted data privacy and access to previously seen data while imposing no constraints on the training overhead. This is unreasonable for FCL applications in real-world scenarios, where edge devices are primarily constrained by resources such as storage, computational budget, and label rate. We revisit this problem with a large-scale benchmark and analyze the performance of state-of-the-art FCL approaches under different resource-constrained settings. Various typical FCL techniques and six datasets in two incremental learning scenarios (Class-IL and Domain-IL) are involved in our experiments. Through extensive experiments amounting to a total of over 1,000+ GPU hours, we find that, under limited resource-constrained settings, existing FCL approaches, with no exception, fail to achieve the expected performance. Our conclusions are consistent in the sensitivity analysis. This suggests that most existing FCL methods are particularly too resource-dependent for real-world deployment. Moreover, we study the performance of typical FCL techniques with resource constraints and shed light on future research directions in FCL.
Index Terms:
Federated Learning, Federated Continual Learning, Resource Constraints.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
Federated Learning (FL) emerges as a distributed learning paradigm, facilitating the collaborative training of a global deep learning model among edge clients while ensuring the privacy of locally stored data[1, 2, 3]. Recently, FL has garnered significant interest and found applications in diverse domains, including recommendation systems [4, 5] and smart healthcare solutions [6, 7].
@@
Typically, FL has been actively studied in a static setting, where the number of training samples does not change over time. However, in a realistic FL application, each client may continue collecting new data and train the local model with streaming tasks, leading to performance degradation on previous tasks [8, 9, 10]. Such phenomenon is known as catastrophic forgetting [11] in the Continual Learning (CL) paradigm. This challenge is further compounded in FL settings, where the local data remains inaccessible to others, and training on clients is constrained by limited resources.
@@
TABLE I: Primary Directions of Progress in FCL. Analysis of three kinds of classifications in the study of the CL system (replay, regularization, and network), with bold highlighting the main contribution. Besides, we further summarize four typical FCL techniques (Sample Caching, Data Synthesis, Knowledge Distillation, and Network Extension) from existing works and focus on two common FCL scenarios. Here ”Constraint” denotes the research based on whether the additional resource overheads are required (Memory Buffer, Computational Budget and Label Rate).
@@
FedAvg [1]
FedProx [12]
Replay TARGET [13] Exemplar Sample
FedCIL [14] Generation & Alignment
Re-Fed [15] Synergistic Replay
AF-FCL [16] Accurate Forgetting
Regularization GLFC [17] Class-Aware Loss
FOT [18] Orthogonality Projection
CFeD [19] Knowledge Distillation
FedET [20] Pre-training Backbone
MFCL [21] Data-Free Distillation
LGA [22] Category-Aware Loss
Network FCL-BL [23] Broad Learning
FedWeIT [24] Parameter Decomposition
pFedDIL [25] Knowledge Matching
To address this issue, researchers have studied federated continual learning (FCL), which enables each client to learn from a local private and incremental task stream continuously. One mainstream technique is to re-train the model with cached or synthetic samples later, a strategy referred to as data replay. FedCIL is proposed in [14] to learn a generative network and reconstruct previous samples for replay, improving the retention of previous information. The authors in [15] propose to selectively cache important samples from each task and train the local model with both cached samples and the samples from the current task. Another promising solution for FCL is to limit the update of model parameters to new tasks, a process known as parameter regularization. The works in [17, 22] concentrate on federated class-incremental learning, addressing scenario-specific challenges through the computation of supplementary class-imbalance losses to train a unified global model. It is studied in [18] that the strategy of projecting the parameters of different tasks onto different orthogonal subspaces prevents new tasks from overwriting previous task parameters. In addition, the authors in [24, 23, 26] extend the backbone model with multi-head classifiers to isolate the task-specific parameters to prevent catastrophic forgetting.
However, the current FCL literature overlooks a key necessity for practical real deployment of such algorithms. In particular, most existing methods focus on offline federated continual learning where, despite limited access to previous data, training algorithms do not have restrictions on the training resources. For example, Snapchat [27], in 2017, reported an influx of over 3.5 billion short videos daily from users across the globe. These videos had to be instantly processed for various tasks, from image rating and recommendation to hate speech and misinformation detection. Otherwise, new streaming data will accumulate until training is completed, causing server delays and worsening user experience. Different from the conditions in the centralized server, distributed devices (usually mobile or edge devices) pose greater challenges to the training resources. In this paper, we mainly focus on the following three mainstream limited resources, which are expensive on distributed devices.
@@
Computational Budget. The computational budgets of continual learning (CL) algorithms tend to exceed those of static data-based methods, as the model must converge on each incremental task prior to learning new ones. Amidst the increasing influx of streaming data, it is paramount for each client to optimize their computational efficiency. However, distributed devices (e.g., IoT and mobile devices) are usually equipped with portable hardware and sacrifice powerful computing resources. A feasible way for distributed devices to improve computational efficiency is to rent cloud platform computing resources, but this is expensive. According to [28], executing a CL algorithm on the CLEAR benchmark [29], which performs 300,000 iterations, incurs approximately a cost of 100\$ on an A100 Google instance, equating to 3\$ per hour for a single GPU. However, in existing methods, such as those based on replay, additional computational budgets are required for retraining the cached samples or training generative models to produce pseudo-data. On the other hand, regularization-based methods may necessitate extra computation for knowledge distillation. In FCL scenarios, limiting the computational budget is necessary for reducing the overall cost.
@@
This raises the question: “Do existing federated continual learning algorithms perform well under resource-constrained settings?” To address this question, we exhaustively study federated continual learning systems, analyzing the effect of the primary directions of progress proposed in the literature in the setting where algorithms are permitted to limit all three training resources. We evaluate and benchmark at scale various existing FCL methods and summarize four typical FCL techniques (Sample Caching, Data Synthesis, Knowledge Distillation, and Network Extension) that are common in the literature. Evaluation is carried out on six large-scale datasets in two scenarios (Class-IL and Domain-IL), amounting to a total of 1,000+ GPU hours under various settings. As shown in Table I, we conclude the existing FCL methods and analyze their main contributions and typical techniques with training resources involved. We summarize our empirical conclusions in three folds:
@@
II Federated Continual Learning with
Resource Constraints
II-A Problem Formulation
Federated Learning. A typical FL problem can be formalized by collaboratively training a global model for $K$ total clients in FL. We consider each client $k$ can only access to his local private dataset $\mathcal{T}_{k}=\{x^{(i)}_{k},y^{(i)}_{k}\}$, where $x^{(i)}_{k}$ is the $i$-th input data sample and $y^{(i)}_{k}\in\{1,2,\cdots,C\}$ is the corresponding label of $x^{(i)}_{k}$ with $C$ classes. We denote the number of data samples in dataset $\mathcal{T}_{k}$ by $|\mathcal{T}_{k}|$. The global dataset is considered as the composition of all local datasets $\mathcal{T}=\{\mathcal{T}_{1},\mathcal{T}_{2},\cdots,\mathcal{T}_{K}\}$, $\mathcal{T}=\sum_{k=1}^{K}\mathcal{T}_{k}$. The objective of the FL learning system is to learn a global model $w$ that minimizes the total empirical loss over the entire dataset $D$:
$\text{where}\ \mathcal{L}_{k}(w)=\frac{1}{|\mathcal{T}_{k}|}\sum_{i=1}^{|\mathcal{T}_{k}|}\mathcal{L}_{CE}(w;x_{i}^{k},y_{i}^{k}),$ (1)
where $\mathcal{L}_{k}(w)$ is the local loss in the $k$-th client and $\mathcal{L}_{CE}$ is the cross-entropy loss function that measures the difference between the prediction and the ground truth labels.
Continual Learning. In the standard CL (non-federated environment), a model learns from a sequence of streaming tasks $\mathcal{T}^{1},\mathcal{T}^{2},\cdots,\mathcal{T}^{n}$} where $\mathcal{T}^{t}$ denotes the $t$-th task of the dataset. Here $\mathcal{T}^{t}=\sum_{i=1}^{N^{t}}(x^{(i)}_{t},y^{(i)}_{t})$, which has $N^{t}$ pairs of sample data $x^{(i)}_{t}\in\mathcal{X}^{t}$ and corresponding label $y^{(i)}_{t}\in\mathcal{Y}^{t}$. We use $\mathcal{X}^{t}$ and $\mathcal{Y}^{t}$ to represent the domain space and label space for the $t$-th task, which has $|\mathcal{Y}^{t}|$ classes and $\mathcal{Y}$ = $\bigcup_{t=1}^{n}\mathcal{Y}^{t}$ where $\mathcal{Y}$ denotes the total classes of all time. Similarly, we use $\mathcal{X}$ = $\bigcup_{t=1}^{n}\mathcal{X}^{t}$ to denote the total domain space for tasks of all time. In this paper, we focus on two types of CL scenarios: (1) Class-Incremental Task: all tasks share the same domain space, i.e., $\mathcal{X}^{1}=\mathcal{X}^{t},\forall t\in[n]$. As the sequence of learning tasks arrives, the number of the classes may change, i.e., $\mathcal{Y}^{1}\neq\mathcal{Y}^{t},\forall t\in[n]$. (2) Domain-Incremental Task: all tasks share the same number of classes i.e., $\mathcal{Y}^{1}=\mathcal{Y}^{t},\forall t\in[n]$. As the sequence of tasks arrives, the client needs to learn the new task while their domain and data distribution changes, i.e., $\mathcal{X}^{1}\neq\mathcal{X}^{t},\forall t\in[n]$.
Federated Continual Learning. We further consider CL in a federated setting. We aim to train a global model for $K$ total clients and assume that client $k$ can only access the local private streaming tasks $\mathcal{T}_{k}^{1},\mathcal{T}_{k}^{2},\cdots,\mathcal{T}_{k}^{n}$}. When the $t$-th task comes, the goal is to train a global model $w^{t}$ overall $t$ tasks $\mathcal{T}^{t}=\{\sum_{n=1}^{t}\sum_{k=1}^{K}\mathcal{T}_{k}^{n}\}$, which can be formulated as :
II-B Dissecting Federated Continual Learning Systems
@@
Federated continual learning methods typically propose a system of multiple components that jointly help improve learning performance. In this section, we analyze FCL systems and dissect them into their underlying techniques. This helps to analyze and isolate the role of different components with different techniques under our resource-constrained settings and helps us to understand the most significant techniques. Overall, we summarize four major techniques from main contributions in Table I on which advances have focused, namely, Sample Caching, Data Synthesis, Knowledge Distillation, and Network Extension.
@@
(1) Sample Caching. Rehearsing samples from previous tasks is a basic approach in FCL. Particularly when access to previous samples is restricted to a small memory, they are used to select which samples from the stream will update the memory. Recent FCL works in [15, 30] propose to synergistically cache samples based on both the local and global understanding. Other methods have not considered the sample selection strategy in FCL. For a fair comparison in our constrained memory buffer setup, we employ random sampling equally to corresponding methods.
@@
(2) Data Synthesis. Considering the security and privacy of data in some scenarios, or data protection under GDPR, the data of previous tasks cannot be cached locally. Some studies use generative models to learn the distribution of previous tasks and generate synthetic data for data replay. The authors in [14, 13] employ the GAN model [31] to generate synthetic samples while [16] explore the NF model [32] to learn the data distribution accurately. In our setting, the synthetic data will also occupy the local memory buffer and the additional training cost brought by the generative model will share the computational budget with the target model.
@@
(3) Knowledge Distillation. As a popular approach, knowledge distillation preserves model performance on previous tasks by distilling the knowledge from the old model (teacher) to the new model (student). In the FCL, the distillation can be done on either the client side or the server side. On the client side, the student model can be the current model, while the teacher model is the one that has been trained for many previous tasks. On the server side, the student model is always the global model and the teacher model can be denoted as the ensemble prediction by each participating client. In this paper, we only focus on the distillation done on the client side since the resources of the server are not so scarce. The distillation will bring an extra computational budget compared with the basic baseline (FedAvg). Moreover, additional distillation data used on the client side will also be concluded in the memory buffer overhead.
@@
(4) Network Extension. Network extension strategies have been used for two objectives. On the one hand, it has been hypothesized that the large difference in the magnitudes of the weights associated with different classes in the last fully connected layer is among the key reasons behind catastrophic forgetting. There has been a family of different methods addressing this problem [24, 20, 18]. On the other hand, several works attempt to adapt the model architecture according to the data. This is done by only training part of the model or by directly expanding the model when data is presented [33, 26]. However, most of the previous techniques in this area do not apply to our setup. Most of this line of work [24, 20, 18], assumes a task-incremental setting, where test samples are known to what set of tasks they belong with known boundaries between streaming tasks. For a pair comparison, we modify these methods to make a correct classification among all classes from all streaming tasks.
III Experiments
@@
  • How do resource constraints affect the performance of existing FCL methods? (Section III-B1)
@@
  • How do typical techniques studied in FCL literature work under different resource-constrained settings? (Section III-B2)
III-A Experiment Setup
III-A1 Datasets.
We conduct our experiments with heterogeneously partitioned datasets over two federated incremental scenarios with six datasets.
Class-Incremental Task Dataset: New classes are incrementally introduced. The dataset starts with a subset of classes, and new classes are added in subsequent stages, allowing models to learn and adapt to an increasing number of classes.
(1) CIFAR-10 [34]: A dataset with 10 object classes, including various everyday objects, animals, and vehicles. It consists of 50,000 training images and 10,000 test images.
(2) CIFAR-100 [34]: Similar to CIFAR-10, but with 100 fine-grained object classes. It has 50,000 training images and 10,000 test images.
(3) Tiny-ImageNet [35]: A subset of the ImageNet dataset with 200 object classes. It contains 100,000 training images, 10,000 validation images, and 10,000 test images.
Domain-Incremental Task Dataset: New domains are introduced gradually. The dataset initially contains samples from a specific domain, and new domains are introduced at later stages, enabling models to generalize to new domains.
(1) Digit-10: Digit-10 dataset contains 10 digit categories in four domains: MNIST [36], EMNIST [37], USPS [38], SVHN [39].Each dataset is a digit image classification dataset of 10 classes in a specific domain, such as handwriting style.
  • MNIST: A dataset of handwritten digits with a training set of 60,000 examples and a test set of 10,000 examples.
  • EMNIST: An extended version of MNIST that includes handwritten characters (letters and digits) with a training set of 240,000 examples and a test set of 40,000 examples.
  • USPS: The United States Postal Service dataset consists of handwritten digits with a training set of 7,291 examples and a test set of 2,007 examples.
  • SVHN: The Street View House Numbers dataset contains images of house numbers captured from Google Street View, with a training set of 73,257 examples and a test set of 26,032 examples.
(2) Office-31 [40]: A dataset with images from three different domains: Amazon, Webcam, and DSLR. It consists of 31 categories, with each domain having around 4,100 images.
(3) Office-Caltech-10 [41]: A dataset with images from four different domains: Amazon, Caltech, Webcam, and DSLR. It consists of 10 object categories, with each domain having around 2,500 images.
TABLE II: Performance comparison of various methods in two incremental scenarios.

### v2

@@ Abstract
Federated Continual Learning (FCL) aims to enable sequential privacy-preserving model training on streams of incoming data that vary in edge devices by preserving previous knowledge while adapting to new data. Current FCL literature focuses on restricted data privacy and access to previously seen data while imposing no constraints on the training overhead. This is unreasonable for FCL applications in real-world scenarios, where edge devices are primarily constrained by resources such as storage, computational budget, and label rate. We revisit this problem with a large-scale benchmark and analyze the performance of state-of-the-art FCL approaches under different resource-constrained settings. Various typical FCL techniques and six datasets in two incremental learning scenarios (Class-IL and Domain-IL) are involved in our experiments. Through extensive experiments amounting to a total of over 1,000+ GPU hours, we find that, under limited resource-constrained settings, existing FCL approaches, with no exception, fail to achieve the expected performance. Our conclusions are consistent in the sensitivity analysis. This suggests that most existing FCL methods are particularly too resource-dependent for real-world deployment. Moreover, we study the performance of typical FCL techniques with resource constraints and shed light on future research directions in FCL.
1 Introduction
@@
Federated Learning (FL) emerges as a distributed learning paradigm, facilitating the collaborative training of a global deep learning model among edge clients while ensuring the privacy of locally stored data mcmahan2017communication ; wang2024fedcda ; wang2023dafkd ; li2025re . Recently, FL has garnered significant interest and found applications in diverse domains, including recommendation systems www_fr1 ; du2020federated ; li2025personalized ; li2025systematic and smart healthcare solutions xu2021federated ; nguyen2022federated ; zhu2024hypernetwork .
@@
Typically, FL has been actively studied in a static setting, where the number of training samples does not change over time. However, in a realistic FL application, each client may continue collecting new data and train the local model with streaming tasks, leading to performance degradation on previous tasks yang2024federated ; li2024unleashing ; lifedssi . Such a phenomenon is known as catastrophic forgetting ganin2016domain in the Continual Learning (CL) paradigm. This challenge is further compounded in FL settings, where the local data remains inaccessible to others, and training on clients is constrained by limited resources.
@@
Table 1: Primary Directions of Progress in FCL. Analysis of three kinds of classifications in the study of the CL system (replay, regularization, and network), with bold highlighting the main contribution. Besides, we further summarize four typical FCL techniques (Sample Caching, Data Synthesis, Knowledge Distillation, and Network Extension) from existing works and focus on two common FCL scenarios. Here "Constraint" denotes the research based on whether the additional resource overheads are required (Memory Buffer, Computational Budget and Label Rate).
@@
FedAvg mcmahan2017communication
FedProx li2020federated
Replay
TARGET zhang2023target Exemplar Sample
FedCIL qi2023better Generation & Alignment
Re-Fed li2024efficient Synergistic Replay
AF-FCL wuerkaixi2023accurate Accurate Forgetting
Regularization
GLFC dong2022federated Class-Aware Loss
FOT bakman2023federated Orthogonality Projection
CFeD ma2022continual Knowledge Distillation
FedET liu2023fedet Pre-training Backbone
MFCL babakniya2024data Data-Free Distillation
LGA dong2023no Category-Aware Loss
Network
FCL-BL le2021federated Broad Learning
FedWeIT yoon2021federated Parameter Decomposition
pFedDIL li2024pfeddil Knowledge Matching
To address this issue, researchers have studied federated continual learning (FCL), which enables each client to learn from a local private and incremental task stream continuously. One mainstream technique is to re-train the model with cached or synthetic samples later, a strategy referred to as data replay. FedCIL is proposed in qi2023better to reconstruct previous samples with a learnable generative network for replay, improving the retention of previous information. The authors in li2024efficient propose to discern important samples from streaming tasks and update the local model with both cached samples and the current task’s samples. Another promising solution for FCL is to limit the update of model parameters to new tasks, a process known as parameter regularization. The works in dong2022federated ; dong2023no concentrate on federated class-incremental learning, addressing scenario-specific challenges through the computation of supplementary class-imbalance losses to train a unified global model. It is studied in bakman2023federated that projecting the different tasks’ parameters onto orthogonal subspaces prevents new tasks from overwriting previous task parameters. In addition, the authors in yoon2021federated ; le2021federated ; yuan2023peer extend the backbone model with multi-head classifiers to isolate the task-specific parameters to prevent forgetting.
However, the current FCL literature overlooks a key necessity for practical real deployment of such algorithms. In particular, most existing methods focus on offline federated continual learning where, despite limited access to previous data, training algorithms do not have restrictions on the training resources. For example, Snapchat zhang2024continual , in 2017, reported an influx of over 3.5 billion short videos daily from users across the globe. These videos had to be instantly processed for various tasks, from image rating and recommendation to hate speech and misinformation detection. Otherwise, new streaming data will accumulate until training is completed, causing server delays and worsening user experience. Different from the conditions in the centralized server, distributed devices (usually mobile or edge devices) pose greater challenges to the training resources. In this paper, we mainly focus on the following three mainstream resources, which are expensive on distributed devices.
@@
Computational Budget. The computational budgets of continual learning (CL) algorithms tend to exceed those of static data-based methods, as the model must converge on each incremental task prior to learning new ones. Amidst the increasing influx of streaming data, it is paramount for each client to optimize their computational efficiency. However, distributed devices (e.g., IoT and mobile devices) are usually equipped with portable hardware and sacrifice powerful computing resources. A feasible way for distributed devices to improve computational efficiency is to rent cloud platform computing resources, but this is expensive. According to prabhu2023computationally , executing a CL algorithm on the CLEAR benchmark lin2022clear , which performs 300,000 iterations, incurs approximately a cost of 100\$ on an A100 Google instance, equating to 3\$ per hour for a single GPU. However, in existing methods, such as those based on replay, additional computational budgets are required for retraining the cached samples or training generative models to produce pseudo-data. On the other hand, regularization-based methods may necessitate extra computation for knowledge distillation. In FCL scenarios, limiting the computational budget is necessary for reducing the overall cost.
@@
This raises the question: “Do existing federated continual learning algorithms perform well under resource-constrained settings?” To address this question, we exhaustively study federated continual learning systems, analyzing the effect of the primary directions of progress proposed in the literature in the setting where algorithms are permitted to limit all three training resources. We evaluate and benchmark at scale various existing FCL methods and summarize four typical FCL techniques (Sample Caching, Data Synthesis, Knowledge Distillation, and Network Extension) that are common in the literature. Evaluation is carried out on six large-scale datasets in two scenarios (Class-IL and Domain-IL), amounting to a total of 1,000+ GPU hours under various settings. As shown in Table 1, we conclude the existing FCL methods and analyze their main contributions and typical techniques with training resources involved. We summarize our empirical conclusions in three folds:
@@
2 Dissecting Federated Continual Learning Systems
@@
Federated continual learning methods typically propose a system of multiple components that jointly help improve learning performance. The problem formulation is illustrated in Appendix B. In this section, we analyze FCL systems and dissect them into their underlying techniques. This helps to analyze and isolate the role of different components with different techniques under our resource-constrained settings and helps us to understand the most significant techniques. Overall, we summarize four major techniques from main contributions in Table 1.
@@
(1) Sample Caching. Rehearsing samples from previous tasks is a basic approach in FCL. Particularly when access to previous samples is restricted to a small memory, they are used to select which samples from the stream will update the memory. Recent FCL works in li2024efficient ; li2024sr propose to synergistically cache samples based on both the local and global understanding. Other methods have not considered the sample selection strategy in FCL. For a fair comparison in our constrained memory buffer setup, we employ random sampling equally to corresponding methods.
@@
(2) Data Synthesis. Considering the security and privacy of data in some scenarios, or data protection under GDPR, the data of previous tasks cannot be cached locally. Some studies use generative models to learn the distribution of previous tasks and generate synthetic data for data replay. The authors in qi2023better ; zhang2023target employ the GAN model pmlr-v70-odena17a to generate synthetic samples while wuerkaixi2023accurate explore the NF model pmlr-v37-rezende15 to learn the data distribution accurately. In our setting, the synthetic data will also occupy the memory buffer and the additional training cost brought by the generative model will share the computational budget with the target model.
@@
(3) Knowledge Distillation. As a popular approach, knowledge distillation preserves model performance on previous tasks by distilling the knowledge from the old model (teacher) to the new model (student). In the FCL, the distillation can be done on either the client side or the server side. On the client side, the student model can be the current model, while the teacher model is the one that has been trained for many previous tasks. On the server side, the student model is always the global model and the teacher model can be denoted as the ensemble prediction by each participating client. In this paper, we only focus on the distillation done on the client side since the resources of the server are not so scarce. The distillation will bring an extra computational budget compared with the basic FedAvg. Moreover, additional distillation data used on the client side will also be included in the memory buffer overhead.
@@
(4) Network Extension. Network extension strategies have been used for two objectives. On the one hand, it has been hypothesized that the large difference in the magnitudes of the weights associated with different classes in the last fully connected layer is among the key reasons behind catastrophic forgetting. There has been a family of different methods addressing this problem yoon2021federated ; liu2023fedet ; bakman2023federated . On the other hand, several works attempt to adapt the model architecture according to the data. This is done by only training part of the model or by directly expanding the model when data is presented Jung2020ContinualLW ; yuan2023peer . However, most of the previous techniques in this area do not apply to our setup. Most of this line of work yoon2021federated ; liu2023fedet ; bakman2023federated , assumes a task-incremental setting, where test samples are known to what set of tasks they belong with known boundaries between streaming tasks. For a fair comparison, we modify these methods to make a correct classification among all classes from streaming tasks.
3 Experiments
@@
  • How do resource constraints affect the performance of existing FCL methods? (Section 3.2.1)
@@
  • How do typical techniques studied in FCL literature work under different resource-constrained settings? (Section 3.2.2)
3.1 Experiment Setup
3.1.1 Datasets
We conduct our experiments with heterogeneous datasets over two typical scenarios: Class-Incremental Learning and Domain-Incremental Learning on six datasets: CIFAR-10 krizhevsky2009learning , CIFAR-100 krizhevsky2009learning , Tiny-ImageNet le2015tiny , Digit-10, Office-31 saenko2010adapting and Office-Caltech-10 zhang2020impact . More details can be found in Appendix B.1.1.
3.1.2 Baselines
In this paper, we follow the same protocols proposed by mcmahan2017communication ; rebuffi2017icarl to set up FIL tasks. We evaluate our resource-constrained settings with twelve baselines: FedAvg mcmahan2017communication , FedProx li2020federated , FL+ERrolnick2019experience , FCIL dong2022federated , FedCIL qi2023better , Target zhang2023target , AF-FCL wuerkaixi2023accurate , Re-Fed li2024efficient , FOT bakman2023federated and FedWeIT yoon2021federated . More details about each baseline can be found in Appendix B.1.2.
3.1.3 Resource Constraints
We analyze three common constrained resources at edge devices: Memory Buffer, Computational Budget, and Label Rate. Here, we abbreviate the memory buffer size as $M$, which means the number of the cached samples in the edge storage, and the computational budget amount as $B$ that denotes the gradient step during the model update; Label Rate as $R$, which represents the ratio of labeled data and unlabeled data. We have adopted different restriction methods for memory buffers and computational budgets for different baselines. The memory buffer is used to cache samples from previous tasks (Re-Fedli2024efficient , FCILdong2022federated ), synthetic data (FedCILqi2023better , Targetzhang2023target , AF-FCLwuerkaixi2023accurate ), and distillation samples (CFeDma2022continual ). The computational budget is shared by the generative model (FedCILqi2023better , AF-FCLwuerkaixi2023accurate ), personalized model (Re-Fedli2024efficient ), knowledge distillation (FOTbakman2023federated , CFeDma2022continual ). The label rate is equally employed to all baselines, which controls the ratio of the labeled data.
3.1.4 Configurations
Unless otherwise mentioned, we use ResNet18 he2016deep as the backbone model for most methods and use the Dirichlet distribution Dir($\alpha$) to distribute local samples to stimulate data heterogeneity for all tasks where a smaller $\alpha$ indicates higher data heterogeneity. Here we report the final accuracy $A(f)$ when the model finishes the training of the last streaming task and the average accuracy $\bar{A}$ across all streaming tasks. Although we try to eliminate the gap between baseline methods other than the core algorithm design in the experiment, there are still some methods that directly depend on the backbone model itself. In our experiments, FedCILqi2023better employs ACGAN for generative replay, and AF-FCLwuerkaixi2023accurate needs to use a Normalized Flow model to generate the feature space for an accurate forgetting. In addition, we try to adopt the ResNet18 to the FOTbakman2023federated but achieve an unstable result. According to the available code in the original manuscript, we use AlexNet as the backbone model for the FOTbakman2023federated . The feasibility of using the ResNet network and achieving excellent performance through simple parameter tuning in the FOT algorithm is still under investigation.
Hardware. Each experiment setting is run twice and we take each run’s final 10 rounds’ accuracy and calculate the average value and standard variance. We use Adam as an optimizer with a linear learning rate schedule. We set the remaining parameters according to the values in the original open-source code, such as the weight of multiple distillation losses. All experiments are run on 8 RTX 4090 GPUs and 8*2 RTX 3090 GPUs.